# SEEA-R1: Tree-Structured Reinforcement Fine-Tuning for Self-Evolving Embodied Agents

### v1

@@ I.7 Visualization of Policy and Reward Model Interactions
J.2 Prompt for the EmbodiedEval Dataset
@@
J.3 Reward Model
Instruction Prompt for Reward Model A conversation between User and Assistant. The user provides the task and the current state, and the Assistant evaluates the state and determine if the current state is success, failure, or continue. Failure means the user cannot achieve success through further actions, while continue means the user can still achieve success through additional actions.  The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think> <answer> Success/Failure/Continue </answer>.

### v2

@@ 11 Put the cups on the tea table onto the dining table 11 Put the shoes on the shoe rack
N.2 Prompt for the EmbodiedEval Dataset
@@
N.3 Prompt for the real-world experiment.
Instruction Prompt for real-world experiment. Interact with a household to solve a task. You are an intelligent agent in a real-world household environment and your target is to perform actions to complete the task goal. At the beginning of your interactions, you will be given the goal to accomplish. For each of your turn, you will be given the observation of the last turn. You should choose from two actions: "Thought" or "Action". If you choose "Thought", you should first think about the current condition and plan for your future actions, and then output your action in this turn. Your output must strictly follow this format:"Thought: your thoughts. Action: your next action"; If you choose "Action", you should directly output the action in this turn. Your output must strictly follow this format:"Action: your next action". The available actions are: 1. look around 2. turn left 3. turn right 4. turn around 5. move forward {number} steps 6. move backward {number} steps 7. go to{recep} 8. take {obj} from{recep} 9. put {obj} in/on{recep} 10. open{recep} 11. close{recep} 12. use {obj} 13. clean {obj} with{recep} 14. heat {obj} with{recep} 15. done where {obj},{recep} and {number} correspond to objects, receptacles and the number. After your each turn, the environment will give you immediate feedback based on which you plan your next few steps. if the envrionment output "Nothing happens", that means the previous action is invalid and you should try more options. Reminder: 1. The action must be chosen from the given available actions. Any actions except provided available actions will be regarded as illegal. 2. Think when necessary, try to act directly more in the process.
N.4 Prompt for the Multi-modal Generative Reward Model.
Instruction Prompt for MGRM A conversation between User and Assistant. The user provides the task and the current state, and the Assistant evaluates the state and determine if the current state is success, failure, or continue. Failure means the user cannot achieve success through further actions, while continue means the user can still achieve success through additional actions. The assistant first thinks about the reasoning process in the mind and then provides the user with the answer. The reasoning process and answer are enclosed within <think> </think> and <answer> </answer> tags, respectively, i.e., <think> reasoning process here </think> <answer> Success/Failure/Continue </answer>.